OpenStudy (anonymous):

Find two unit vectors orthogonal (perpendicular) to both the given vectors:

OpenStudy (anonymous):

<1, -1, 1> <0, 4, 4>

OpenStudy (abb0t):

find the det.

OpenStudy (amistre64):

i was thinking a cross product meself

OpenStudy (amistre64):

ahh, "determinant" ... yeah

OpenStudy (abb0t):

oh yeah, cross product. Lol. if it equals 1, it's orthogonal.

OpenStudy (anonymous):

My work: Found determent of following matrix: i j k 1 -1 1 0 4 4 \[=-8i -4j + 4k\] and: i j k 0 4 4 1 -1 1 \[=8i + 4j - 4k\] So, translate to unit vectors you should have: <-1, -1, 1> and <1, 1, -1>

OpenStudy (amistre64):

8,4,4 doesnt reduce to 1,1,1

OpenStudy (anonymous):

unit vector

OpenStudy (amistre64):

and 1,1,1 is sqrt3 long, not 1 long

OpenStudy (anonymous):

ah, that's probably what I'm doing wrong..

OpenStudy (amistre64):

<1, -1, 1> <0, 4, 4> x 0 1 y 4 -1 z 4 1 x = (4--4) y = -(0-4) z = (0-4) (8,4,-4) is normal to both ... we can reduce by a factor of 4 (2,1,-1) , now divide it by its length

OpenStudy (amistre64):

+- (2,1,-1) ---------- sqrt6

OpenStudy (anonymous):

thanks